i can't under stand if 3x+2y = -17
2x-5y = 14 the what is the process i can't under stand
what i need is Ex:
this i understand
X+2y=5
y=2x
X+2y=5
x+2(2x)= 5
x+4x=5
x+ 4= 5
5x and 5 divided by 5 y = 1 that i under stand but below i don't get it

THIS!!

Explanation:

Solving:
3x+2y = -17
2x-5y = 14

Using Substitution Method.
Expressing x as a function of y (from the first equation) x = (-17 -2y)/3
Substituting this expression into the second equation:
2(-17+2y)/3-5y = 14
Simplifying this expression:
(-19)y = 76
or
y = -4
Now, substituting y = -4 back into the x expressed as a function of y:
x = (-17-2y)/3 = -3


Answer by rfer(16322)

3x+2y=-17
2x-5y=14
2x=5y+14
x=2.5y+7
--------------
3*(2.5y+7)+2y=-17
7.5y+21+2y=-17
9.5y=-38
y=-4
-----------------
3x+2*-4=-17
3x-8=-17
3x=-9
x=-3
---------------
2*-3-5*-4=14
-6+20=14